# Scalable Composite Likelihood Estimation of Probit Models with Crossed Random Effects

### v2

@@ 6 Bibliographic Notes
7 Conclusions
We have developed a method for crossed random effects for binary regression. It is a new version of composite likelihood that, by exploiting the connection between the Gaussian latent variables and the Gaussian random effects in the probit model, makes for very convenient factors in the composite likelihood.
In simulations, we find that the computational cost of our ARC proposal is $\mathcal{O}(N)$ and its accuracy is competitive with and even more robust than a computationally infeasible Laplace approximation that has a cost which grow superlinearly with $N$. The mean squared errors of our method are very nearly $\mathcal{O}(\min\{R,C\}^{-1})$ for the intercept and $\mathcal{O}(N^{-1})$ for the other regression parameters. They are nearly $\mathcal{O}(R^{-1})$ and $\mathcal{O}(C^{-1})$ for the row- and column-random effect variance parameters, respectively.
In the application to the e-commerce Stitch Fix data, a naive probit model finds that the coefficients of a model with 12 regression coefficients are all very statistically significant. After properly accounting for correlations due to random effects, we find that only eight of them remain significant at the 5% level. The naive analysis might lead to one following false leads. It can even lead one to confidently attribute incorrect signs to some parameters, even though $\mathrm{sign}(\gamma_{j})=\mathrm{sign}(\beta_{j})$. This can happen when $\mathrm{sign}(\hat{\gamma}_{j})\neq\mathrm{sign}(\gamma_{j})$ and the estimate of $\mathrm{var}(\hat{\gamma}_{j})$ is unrealistically small.
One surprise in this work was the difficulty of getting good approximations to even the one dimensional Gaussian integrals that arise. We are grateful for the insights in Ogden (2021) and Shun and McCullagh (1995). We surmise that the difficulties in multidimensional integration for random effects models on big data sets must be even more severe than the univariate ones we saw.
We have used a probit model instead of a logistic one because a Gaussian latent variable is a very natural counterpart to the Gaussian random effects that are the default in random effects models. That connection simplified our modeling and computation. As noted in the introduction, logit and probit link functions are nearly proportional outside of tail regions.
Gibbons and Hedeker (1997) remark that “As in the case of fixed effect models, selection of probit versus logistic response functions appears to have more to do with custom or practice within a particular discipline than differences in statistical properties.” Thus, while a logistic regression approach is outside the scope of the present paper we think that an extension to the logistic case is worth pursuing and that our present findings provide a good starting point. If we had used a logistic instead of the probit model, then the all-likelihood would have required us to approximate further $N$ univariate integrals with respect to a convolution of the Gaussian and logistic probability density functions. That would be much more expensive than the probit model costs, although still within the $\mathcal{O}(N)$ constraint.
Another approach to extending the ARC method to logistic regression uses the fact that the logistic distribution is a scale mixture of the standard normal distribution (Stefanski, 1991). This property is often used in Bayesian computations, see, for example, Narisetty et al. (2018).
@@ 7 Conclusions
The supplementary material contains proofs of the theorems of Section 3, additional simulation results discussed in Section 4, further references to the literature and a summary table of the binary regression standard errors for the Stitch Fix data.
R code for replicating our results is available in the public repository https://github.com/rugbel/arcProbit.
@@ Acknowledgements
We thank Bradley Klingenberg and Stitch Fix for making this data set available. This work was supported in part by grants IIS-1837931 and DMS-2152780 from the U.S. National Science Foundation, IRIDE from Ca’ Foscari University of Venice, and by the research project Latent Variable Models for Complex Data founded by the European Union - NextGenerationEU (MUR DM funds 737/2021).

### v4

@@ 5 Discussion
The supplementary material contains proofs of the theorems of §2, additional simulation results discussed in §3, further references to the literature, as well as a table and two plots about the Stitch Fix application mentioned in §4. The supplement is at
https://artowen.su.domains/reports/ARC-final-supplement.pdf.
R code for replicating our results is available in the public repository
https://github.com/rugbel/arcProbit.
@@ Acknowledgements
We are grateful to Silvia Bianconcini, Blair Bilodeau, Nancy Reid, Alex Stringer, Yanbo Tang and Matt Wand for helpful discussion about generalized linear mixed models, Laplace approximation and adaptive Gauss-Hermite quadrature. We thank Bradley Klingenberg and Stitch Fix for making this data set available. This work was supported in part by grants IIS-1837931 and DMS-2152780 from the U.S. National Science Foundation, IRIDE from Ca’ Foscari University of Venice, and by the research project Latent Variable Models for Complex Data founded by the European Union - NextGenerationEU (MUR DM funds 737/2021).